# On diffusion-based generative models and their error bounds: The log-concave case with full convergence estimates

### v3

@@ Abstract
We provide full theoretical guarantees for the convergence behaviour of diffusion-based generative models under the assumption of strongly log-concave data distributions while our approximating class of functions used for score estimation is made of Lipschitz continuous functions. We demonstrate via a motivating example, sampling from a Gaussian distribution with unknown mean, the powerfulness of our approach. In this case, explicit estimates are provided for the associated optimization problem, i.e. score approximation, while these are combined with the corresponding sampling estimates. As a result, we obtain the best known upper bound estimates in terms of key quantities of interest, such as the dimension and rates of convergence, for the Wasserstein-2 distance between the data distribution (Gaussian with unknown mean) and our sampling algorithm.
@@ 1 Introduction
Our approach introduces a novel auxiliary process that relies solely on known information and uses the density of the solution of this process in the $L^{2}$-accurate score estimation. To further highlight the powerfulness of our approach, we present a motivating example on the case of sampling from a Gaussian distribution with unknown mean. Full theoretical estimates for the convergence properties of the SGM are provided in Wasserstein-2 distance, while our choice of stochastic optimizer for the score approximation is a simple Langevin-based algorithm. Our estimates are explicit and contain the best known dependencies in terms of dimension and rate of convergence (Theorem 3.1). To the best of the authors’ knowledge, these are the first such explicit results. The proofs of these estimates are derived using standard techniques from stochastic calculus.
@@ 3.1 A Motivating Example: Full Estimates for Multivariate Gaussian Initial Data with Unknown Mean
In this section, we consider the case where the data distribution follows a multivariate normal distribution with unknown mean and identity covariance, i.e., $X_{0}\sim\pi_{\mathsf{D}}=\mathscr{N}(\mu,I_{d})$ for some unknown $\mu\in\mathbb{R}^{d}$ with $M=d$. We show that, by using diffusion models, we are able to generate new data from an approximate distribution that is close to $\pi_{\mathsf{D}}$. More precisely, we provide a non-asymptotic convergence bound with explicit constants in Wasserstein-2 distance between the law of the diffusion model and $\pi_{\mathsf{D}}$, which can be made arbitrarily small by appropriately choosing key parameters on the upper bound.
@@
The result in Theorem 3.1 provides theoretical justification for the empirical success of the diffusion models. Crucially, to the best of the authors’ knowledge, this is the first such result which provides a convergence bound with transparent dependence on the parameters involved in the sampling and optimization procedures of the diffusion models.
More concretely, our result is obtained by connecting the diffusion models with the theoretical guarantees of machine learning optimizers. In the optimisation procedure, we use SGLD (15) to solve (6). Since the stochastic gradient $H$ in (14) is strongly convex by Proposition B.1, it has been proved, for instance in (Barkhagen et al., 2021), that, for large enough $\beta>0$, the output of SGLD is an almost minimizer of (6) when $n$ is large. Thus, in the diffusion model, we set $\hat{\theta}=\theta_{n}^{\lambda}$ indicating $\nabla\log p_{t}(x)\approx s(t,\hat{\theta},x)=s(t,\theta_{n}^{\lambda},x)$ for large values of $n$ and for all $t$ and $x$. Crucially, this allows us to use the established convergence results for SGLD to deduce a sampling upper bound for $W_{2}(\mathscr{L}(\widehat{Y}_{K+1}^{\text{EM}}),\pi_{\mathsf{D}})$ with explicit constants. Consequently, this bound can be controlled by any given precision level $\delta>0$ by appropriately choosing $T,\beta,\lambda,n$ and $\gamma$.
Remark 3.3.
The result in Theorem 3.1 achieves the optimal rate of convergence for Euler or Milstein schemes of SDEs with constant diffusion coefficients, i.e. $\gamma$. Furthermore, one notes that the dependence of the dimension on the upper bound in (18) is in the form of $\sqrt{d}$.
Remark 3.4.
@@ 3.2.2 Full Estimates for the General Case
where $C_{1}$, $C_{2}$, $C_{3}$ and $C_{4}$ are given explicitly in Table 2. In addition, the result in (21) implies that for any $\delta>0$, if we choose $0\leq\epsilon<\epsilon_{\delta}$, $T>T_{\delta}$, $0<\varepsilon_{\text{SN}}<\varepsilon_{\text{SN},\delta}$ and $0<\gamma<\gamma_{\delta}$ with $\epsilon_{\delta}$, $T_{\delta}$, $\varepsilon_{\text{SN},\delta}$, and $\gamma_{\delta}$ given in Table 2, then
@@
The error bounds in (21) are not as good as the ones provided in Theorem 3.1 due to the general form of the approximating function $s$. We emphasize that the optimal rate of convergence for the Euler or Milstein scheme of SDEs with constant diffusion coefficients is achieved, i.e. $\gamma^{\alpha}$, using the Lipschitz continuity on the derivative of $s$ in Assumption 3.
Constant Dependency Full Expression
$C_{1}$ $O(\sqrt{M})$ $2(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{M})$
Table 2: Explicit expressions for the constants in Theorem 3.12.
@@ 4.1 Score Approximation Assumptions
We note that, unlike (22) and (23), the expectation in (25) is taken with respect to the algorithm $Y_{k}^{\text{EI}}$ given by the exponential integrator (EI) discretization scheme [^1] This analysis can be extended to the Euler-Maruyama numerical scheme in $Y_{k}^{\text{EM}}$ (10)., which has known density. However, the bounds of (Bortoli, 2022, Theorem H.1) in Wasserstein distance of order one derived under this assumption scale exponentially in problem parameters such as the diameter of the manifold $\mathfrak{M}$ and the inverse of the early stopping time parameter $\epsilon$, i.e. $\exp(O(\text{diam}(\mathfrak{M})^{2}/\epsilon))$. Furthermore, an assumption similar to (25) is considered in a very recent and concurrent result (Gao et al., 2023):
@@ 4.2 Assumptions on the Data Distribution
The vast majority of the results available in the literature are in KL divergence and TV distance. For two general data distributions $\mu$ and $\nu$, there is no known relationship between their KL divergence and their $W_{2}$. Hence, the results available in the literature on the KL divergence cannot be compared easily with the results in 2-Wasserstein distance, and vice versa. Assuming that the data distribution is bounded, some results in (Chen et al., 2023d, a) control the 2-Wasserstein distance by the TV distance. We point out that we do not follow this strategy in our proofs of Theorem 3.1 and Theorem 3.12 and avoid to assume boundedness on the data distribution.
Some convergence results in different metrics can be deduced under the following types of assumptions. Convergence bounds in Wasserstein distance of order one with exponential complexity has been obtained in (Bortoli, 2022) under the so-called manifold hypothesis, namely assuming that the target distribution is supported on a lower-dimensional manifold or is given by some empirical distribution. Moreover, the results in TV distance in (Lee et al., 2022) and in KL divergence (Yang and Wibisono, 2023) assumed that the data distribution satisfies a logarithmic Sobolev inequality and the score function is Lipschitz resulting in convergence bounds characterized by polynomial complexity. By replacing the requirement that the data distribution satisfies a functional inequality with the assumption that $\pi_{\text{D}}$ has finite KL divergence with respect to the standard Gaussian and by assuming that the score function for the forward process is Lipschitz, the authors in (Chen et al., 2023d) managed to derive bounds in TV distance which scale polynomially in all the problem parameters. By requiring only the Lipschitzness of the score at the initial time instead of the whole trajectory, the authors in (Chen et al., 2023a, Theorem 2.5) managed to show, using an exponentially decreasing then linear step size, convergence bounds in KL divergence with quadratic dimensional dependence and logarithmic complexity in the Lipschitz constant. In the work by (Benton et al., 2024), the authors provide convergence bounds in KL divergence that are linear in the data dimension up to logarithmic factors assuming finite second moments of the data distribution.

### v4

@@ Abstract
We provide full theoretical guarantees for the convergence behaviour of diffusion-based generative models under the assumption of strongly log-concave data distributions while our approximating class of functions used for score estimation is made of Lipschitz continuous functions avoiding any Lipschitzness assumption on the score function. We demonstrate via a motivating example, sampling from a Gaussian distribution with unknown mean, the powerfulness of our approach. In this case, explicit estimates are provided for the associated optimization problem, i.e. score approximation, while these are combined with the corresponding sampling estimates. As a result, we obtain the best known upper bound estimates in terms of key quantities of interest, such as the dimension and rates of convergence, for the Wasserstein-2 distance between the data distribution (Gaussian with unknown mean) and our sampling algorithm.
@@ 1 Introduction
Our approach introduces a novel auxiliary process that relies solely on known information and uses the density of the solution of this process in the $L^{2}$-accurate score estimation. To further highlight the powerfulness of our approach, we present a motivating example on the case of sampling from a Gaussian distribution with unknown mean. Full theoretical estimates for the convergence properties of the SGM are provided in Wasserstein-2 distance, while our choice of stochastic optimizer for the score approximation is a simple Langevin-based algorithm. Our estimates are explicit and contain the best known optimal dependencies in terms of dimension and rate of convergence (Theorem 3.1). To the best of the authors’ knowledge, these are the first such explicit results with transparent dependence on the parameters involved in the sampling and optimization combined procedures of the diffusion models. By connecting the diffusion models with the theoretical guarantees of machine learning optimizers via standard stochastic calculus tools, the results in Theorem 3.1, together with the bounds achieved in the more general setting in Theorem 3.10, provide the theoretical justification for the empirical success of the diffusion models.
@@ 3.1 A Motivating Example: Full Estimates for Multivariate Gaussian Initial Data with Unknown Mean
In this section, we consider the case where the data distribution follows a multivariate normal distribution with unknown mean and identity covariance, i.e., $X_{0}\sim\pi_{\mathsf{D}}=\mathscr{N}(\mu,I_{d})$ for some unknown $\mu\in\mathbb{R}^{d}$ with $M=d$. We show that, by using diffusion models, we are able to generate new data from an approximate distribution that is close to $\pi_{\mathsf{D}}$. More precisely, we provide a non-asymptotic convergence bound with explicit constants in Wasserstein-2 distance between the law of the diffusion model and $\pi_{\mathsf{D}}$, which can be made arbitrarily small by appropriately choosing key parameters on the upper bound. In this example, the estimation of the score function reduces to the estimation of the unknown mean by the methods of convex optimization with optimal dependence of the dimension combined with the most efficient sampling method for high-dimensional Gaussian data.
@@
The result in Theorem 3.1 achieves the optimal rate of convergence of order one for Euler or Milstein schemes of SDEs with constant diffusion coefficients. Furthermore, one notes that the dependence of the dimension on the upper bound in (18) is in the form of $\sqrt{d}$. To the best of the authors’ knowledge, the result in Theorem 3.1 is the first convergence bound where the parameters involved in the sampling and optimization steps of the diffusion models appear explicitly. In the optimization procedure, we use SGLD algorithm (15) to solve the problem (6). Since the stochastic gradient $H$ in (14) is strongly convex by Proposition B.1, it has been proved, for instance in (Barkhagen et al., 2021), that, for large enough $\beta>0$, the output of SGLD is an almost minimizer of (6) when $n$ is large. Thus, in the diffusion model, we set $\hat{\theta}=\theta_{n}^{\lambda}$ indicating $\nabla\log p_{t}(x)\approx s(t,\hat{\theta},x)=s(t,\theta_{n}^{\lambda},x)$ for large values of $n$ and for all $t$ and $x$. Crucially, this allows us to use the established convergence results for SGLD to deduce a sampling upper bound for $W_{2}(\mathscr{L}(\widehat{Y}_{K+1}^{\text{EM}}),\pi_{\mathsf{D}})$ with explicit constants in (18). Consequently, this bound can be controlled by any given precision level $\delta>0$ by appropriately choosing $T,\beta,\lambda,n$ and $\gamma$.
@@ 3.2.2 Full Estimates for the General Case
where $C_{1}$, $C_{2}$, $C_{3}$ and $C_{4}$ are given explicitly in Table 3 (Appendix D). In addition, the result in (21) implies that for any $\delta>0$, if we choose $0\leq\epsilon<\epsilon_{\delta}$, $T>T_{\delta}$, $0<\varepsilon_{\text{SN}}<\varepsilon_{\text{SN},\delta}$ and $0<\gamma<\gamma_{\delta}$ with $\epsilon_{\delta}$, $T_{\delta}$, $\varepsilon_{\text{SN},\delta}$, and $\gamma_{\delta}$ given in Table 3, then
@@
The error bounds in (21) are not as good as the ones provided in Theorem 3.1 due to the general form of the approximating function $s$. We emphasize that the optimal rate of convergence of order $\alpha\in\left[\frac{1}{2},1\right]$ for the Euler or Milstein scheme of SDEs with constant diffusion coefficients is achieved using the Lipschitz continuity on the derivative of $s$ in Assumption 3. In the explicit expression of $C_{4}$ in Table 3, the dependence of the dimension is $O(M)$ due to numerical techniques from Kumar and Sabanis (2019) used in the proof of Theorem 3.10 to achieve the optimal rate of convergence.
@@ 4.1 Score Approximation Assumptions
We note that, unlike (22) and (23), the expectation in (25) is taken with respect to the algorithm $Y_{k}^{\text{EI}}$ given by the exponential integrator (EI) discretization scheme [^1] This analysis can be extended to the Euler-Maruyama numerical scheme in $Y_{k}^{\text{EM}}$ (10)., which has known density. However, the bounds of (Bortoli, 2022, Theorem H.1) in Wasserstein distance of order one derived under this assumption scale exponentially in problem parameters such as the diameter of the manifold $\mathfrak{M}$ and the inverse of the early stopping time parameter $\epsilon$, i.e. $\exp(O(\text{diam}(\mathfrak{M})^{2}/\epsilon))$. Furthermore, an assumption similar to (25) is considered in a very recent and concurrent result [^2] The concurrent paper (Gao et al., 2023) appeared few days earlier when the first draft of this work was made available online. (Gao et al., 2023):
@@ 4.2 Assumptions on the Data Distribution
The vast majority of the results available in the literature are in KL divergence and TV distance. For two general data distributions $\mu$ and $\nu$, there is no known relationship between their KL divergence and their $W_{2}$. However, for strongly log-concave data distributions, a bound on the Wasserstein distance of order two in terms of KL divergence follows from an extension of Talagrand’s inequality (Gozlan and Léonard, 2010, Corollary 7.2).
Some convergence results in different metrics can be deduced under the following types of assumptions. Convergence bounds in Wasserstein distance of order one with exponential complexity has been obtained in (Bortoli, 2022) under the so-called manifold hypothesis, namely assuming that the target distribution is supported on a lower-dimensional manifold or is given by some empirical distribution. Moreover, the results in TV distance in (Lee et al., 2022) and in KL divergence (Yang and Wibisono, 2023) assumed that the data distribution satisfies a logarithmic Sobolev inequality and the score function is Lipschitz resulting in convergence bounds characterized by polynomial complexity. By replacing the requirement that the data distribution satisfies a functional inequality with the assumption that $\pi_{\text{D}}$ has finite KL divergence with respect to the standard Gaussian and by assuming that the score function for the forward process is Lipschitz, the authors in (Chen et al., 2023d) managed to derive bounds in TV distance which scale polynomially in all the problem parameters. By requiring only the Lipschitzness of the score at the initial time instead of the whole trajectory, the authors in (Chen et al., 2023a, Theorem 2.5) managed to show, using an exponentially decreasing then linear step size, convergence bounds in KL divergence with quadratic dimensional dependence and logarithmic complexity in the Lipschitz constant. In the work by (Benton et al., 2024), the authors provide convergence bounds in KL divergence that are linear in the data dimension up to logarithmic factors.
We summarise the results of Chen et al. (2023d, a); Benton et al. (2024) and compare them to ours in Table 2. A careful examination of (Chen et al., 2023a, Proof of Theorem 2.2) and (Benton et al., 2024, Proof of Theorem 1 and Corollary 1) reveals that the authors require to have uniqueness of solutions for the backward SDE (3). Therefore, additional integrability conditions on the score function depending on the type of uniqueness of solutions considered are still needed.
We emphasize that in Theorem 3.10, we do not assume the score function to be Lipschitz continuous with a uniformly bounded Lipschitz constant. This is particularly useful for future work in nonconvex settings, where the upper bound estimates will be independent of the (potentially large) Lipschitz constant of the score function, which could otherwise hide additional dimensional dependencies. The requirement $\int_{\epsilon}^{T}|\nabla\log p_{t}(0)|^{2}\text{d}t<\infty$ in Assumption 2 is weaker than the Lipschiz assumption on the score function, but it is still difficult to verify in practical applications.
@@
Optimization  estimate Regularity  condition Metric Complexity Result
None $\forall t$, $\nabla\log p_{t}$, $L$-Lipschitz $\text{TV}(\mathscr{L}(\widehat{Y}_{K+1}),\pi_{\text{D}})$ $\tilde{O}\left(\frac{ML^{2}}{\delta^{2}}\right)$ (Chen et al., 2023d, Theorem 2)
None $\forall t$, $\nabla\log p_{t}$ $L$-Lipschitz $\sqrt{\text{KL}(\mathscr{L}(\widehat{Y}_{K+1}^{\text{EI}})||\pi_{\text{D}})}$ $\tilde{O}\left(\frac{ML^{2}}{\delta^{2}}\right)$ (Chen et al., 2023a, Theorem 2.1)
None Integrability conditions on $\nabla\log p_{t}$ for the uniqueness of $(Y_{t})_{t\in[0,T]}$ $\sqrt{\text{KL}(\mathscr{L}(\widehat{Y}_{K}^{\text{EI}})||\pi_{\text{D}}^{\underline{\epsilon}})}$ $\tilde{O}\left(\frac{M^{2}\log^{2}(1/\epsilon)}{\delta^{2}}\right)$ (Chen et al., 2023a, Theorem 2.2)
None Integrability conditions on $\nabla\log p_{t}$ for the uniqueness of $(Y_{t})_{t\in[0,T]}$ $\sqrt{\text{KL}(\mathscr{L}(\widehat{Y}_{K}^{\text{EI}})||\pi_{\text{D}}^{\underline{\epsilon}})}$ $\tilde{O}\left(\frac{M\log^{2}(1/\epsilon)}{\delta^{2}}\right)$ (Benton et al., 2024, Corollary 1)
Yes $\pi_{\text{D}}\sim\mathscr{N}(\mu,I_{M})$ $W_{2}(\mathscr{L}(\widehat{Y}_{K+1}^{\text{EM}}),\pi_{\mathsf{D}})$ $\tilde{O}(\sqrt{M}\log(\delta^{-1}))$ Theorem 3.1
Yes Assumption 2 $W_{2}(\mathscr{L}(\widehat{Y}_{K}^{\text{EM}}),\pi_{\mathsf{D}})$ $\tilde{O}(M^{1/\alpha}(\widetilde{\varepsilon}_{\text{AL}}\epsilon)^{1/(2\alpha)}\log(\delta^{-1}))$ Theorem 3.10
Table 2: Summary of previous bounds and our results. Bounds expressed in terms of the number of steps required to guarantee an error of at most $\delta$ in the stated metric. The variance -$\underline{\epsilon}$ Gaussian perturbation of the target data distribution is denoted by $\pi_{\text{D}}^{\underline{\epsilon}}$. All the bounds assume that $\pi_{\text{D}}$ has finite second moments.
@@ C.3 Proof of the Preliminary Results for the General Case
Appendix D Table of Constants
Table 3 displays full expressions for constants which appear in Theorem 3.10.
Constant Dependency Full Expression
$C_{1}$ $O(\sqrt{M})$ $2(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{M})$
Table 3: Explicit expressions for the constants in Theorem 3.10.